# Inside you are many wolves: Using cognitive models to interpret value trade-offs in LLMs

### v1

@@ 8 Conclusion
The internal mechanisms of large language models are often opaque to external observers. Yet, understanding the extent to which their internal trade-offs resemble our own is important to their success as agents, assistants, and judges, and our ability to shape their training towards our desired visions of these applications. The present work continues the fruitful line of research in computational cognitive science that seeks to model human value-trade-offs [84, 39, 69, 13, 70], and connects it to the complementary goals of Inverse Reinforcement Learning. We propose using a cognitively interpretable model of pragmatic language use as a means of understanding LLMs’ value trade-offs as a result of reasoning and alignment. We believe this tool provides a valuable mechanism for guiding model development–enabling the formation of fine-grained hypotheses about high-level behavioral concepts, understanding the extent of training needed to achieve desired model values, and shaping recipes for higher-order reasoning and alignment.
@@ Acknowledgements
We thank the members of the Harvard Computation, Cognition, and Development Lab, as well as Ekdeep Singh Lubana and Hidenori Tanaka for their for their helpful comments and discussion. This material is based upon work supported by the NSF Graduate Research Fellowship under Grant No. DGE 2140743 to SKM. RZ is supported by a Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199. RZ and SKM are supported by Kempner Institute Graduate Research Fellowships. TU was supported by the Jacobs Foundation.
@@
Appendix A Experimental details
A.1 Data
The original experimental vignettes from Yoon et al., [92] can be found here.
A.2 Evaluating LLM responses
@@
To handle LLM generations that did not adhere to the specified multiple-choice format, we used the gpt-4o-2024-08-06 checkpoint of GPT-4o as a judge prompted with the following:
@@
A.3 Manipulations of vignette framing
For our main experimental task, in addition to the third person framing of the vignettes as they were presented in Yoon et al., [92], we test manipulations of framing to simulate the LLMs’ use as an assistant (first person) and agent (second person). For each case, the following expression of the speaker’s true opinion was appended to the scenario as described in the main text, with the relevant framing of the final model query (replacing [speaker] with the appropriate conjugations of “I” and “you”, respectively):
@@ A.3 Manipulations of vignette framing
Appendix B Implementation details
B.1 Open-source model training
For our open source model suite training runs (Section 4), we provide hyperparameter details in Table 2. We use an internal cluster of 80GB H100 GPUs to conduct SFT, DPO, and PPO training runs. For DPO and SFT, training can be done on 4 H100 GPUs with gradient accumulation, with training for 1 epoch taking 3 hours and 6 hours for UltraFeedback and Anthropic HH-RLHF respectively. For PPO, we use 8 H100 GPUs taking 6 hours and 16 hours for UltraFeedback and Anthropic HH-RLHF respectively.
@@ B.2 Cognitive model
We generally follow the modeling assumptions described in Yoon et al., [92], with one exception: where the original model assumes that negated expressions such as “not amazing” have more words and are thus slightly more costly for people to produce, we omit this additional cost and assume that each of the eight utterances are equally costly for an LLM.
Parameter values
The input to the sampling-based inference algorithm, $\mathcal{M}$, was count data transformed proportionally from an LLM’s averaged utterance preferences across vignettes and random combinations of names. For each true state $s$, we mapped an LLM’s utterance distribution $P_{\mathcal{LLM}}(u|s)$ to frequency counts by a scaling factor of total count $|\mathcal{M}|$. We set the total count as 130 (10 name combinations $\times$ 13 vigenttes) for each true state. For example, under the true state “1 star”, if an LLM’s response in the utterance preference task assigns a normalized probability of 0.323 to the utterance “not good” out of the eight possible utterance options, then the corresponding count data $\mathcal{M}_{\text{1 star, ``not good''}}$ for “not good” under the state of “1 star” would be the rounded number of $0.323\times 130\approx 42$.
Appendix C Intermediate results
C.1 Distribution of LLMs’ responses on polite speech task
@@
Figures 4 through 13 show the raw distributions of LLMs’ responses on the main polite speech task for each of the 5 possible true states (1 to 5 stars) in our experimental vignettes. Each figure shows the results for a particular alignment method (DPO or PPO), wherein rows correspond to various combinations of base model and feedback dataset, and columns correspond to vignette framing.
@@
C.2 Literal semantics sub-task
@@
C.3 Fitting LLMs’ responses to first-order speaker model $S_{1}$
Closed-source model suite
To verify the viability of the parameter values inferred by our complete $S_{2}$ speaker model, we test a simpler version of the cognitive model that exits at $S_{1}$, the first-order speaker within $S_{2}$. Figure 16 shows these results for the closed-source model suite. The inferred values of the parameter $\phi$ from this model, roughly match those of the second-order speaker model’s $\phi$.
Figure 16: Inferred values of $\phi$ for simplified first-order speaker model $S_{1}$ for the closed-source model suite.

### v3

@@ 7 Conclusion
The internal mechanisms of large language models are often opaque to external observers. Yet, understanding the extent to which their internal trade-offs resemble our own is important to their success as agents, assistants, and judges, and our ability to shape their training towards our desired visions of these applications. The present work continues the fruitful line of research in computational cognitive science that seeks to model human value-trade-offs (Ullman et al., 2009; Jern & Kemp, 2014; Powell, 2022; Davis et al., 2023; Qian et al., 2024), and connects it to the complementary goals of Inverse Reinforcement Learning. We propose using a cognitively interpretable model of pragmatic language use as a means of understanding LLMs’ value trade-offs as a result of reasoning and alignment. While our work studies a particular set of value trade-offs, we show that this method is responsive to diverse aspects of a rapidly evolving LLM landscape. We believe this tool provides a valuable mechanism for guiding model development–enabling the formation of fine-grained hypotheses about high-level behavioral concepts, understanding the extent of training needed to achieve desired model values, and shaping recipes for higher-order reasoning and alignment.
@@ Acknowledgements
We thank the members of the Harvard Computation, Cognition, and Development Lab, especially Felix Sosa and Eric Bigelow, as well as Ekdeep Singh Lubana and Hidenori Tanaka for their for their helpful comments and discussion. This material is based upon work supported by the NSF Graduate Research Fellowship under Grant No. DGE 2140743 to SKM. RZ is supported by a Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199. RZ and SKM are supported by Kempner Institute Graduate Research Fellowships. TU was supported by the Jacobs Foundation.
@@
Appendix A Background
A.1 Value alignment in LLMs
A substantial body of work on aligning large language models (LLMs) has focused on optimizing models to reflect human preferences. Reinforcement learning-based methods—such as Reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022a) and Reinforcement Learning from AI Feedback (RLAIF) (Bai et al., 2022b)—as well as offline preference optimization techniques like Direct Preference Optimization (DPO) and variants (Rafailov et al., 2023; Ethayarajh et al., 2024; Hong et al., 2024; Park et al., 2024b), have become standard components of the LLM alignment pipeline. These methods are widely believed to underlie many of the human-like behaviors exhibited by current models (Ji et al., 2024). While off-policy methods and the use of static datasets are more efficient and easy to implement, prior work has shown that online methods are superior for preference learning (Tajwar et al., 2024; Tang et al., 2024; Xu et al., 2024). However, prior work has also shown that the resulting models after preference fine-tuning generally show a lack of linguistic and conceptual diversity, which suggests a difficulty in maintaining multiplicity (Kirk et al., 2024; janus, 2022; Padmakumar & He, 2024; Park et al., 2024a; O’Mahony et al., 2024; Murthy et al., 2025; West & Potts, 2025).
Recently, reinforcement learning-based finetuning has become popular for improving mathematical reasoning and coding abilities in models, where rewards are verifiable as opposed to coming from a learned reward model (Zelikman et al., 2022; Lambert et al., 2024; Jaech et al., 2024; Guo et al., 2025; Shao et al., 2024; Team et al., 2025). Such ‘reasoning models’ exhibit certain characteristics such as having longer and more expressive chains of thought (Wei et al., 2022). However, it is unclear what model behavior is elicited— even unintentionally— as a result of optimizing the verifiable rewards in these constricted domains; for instance, DeepSeek R1 underwent an additional stage of preference finetuning for safety alignment (Guo et al., 2025). In spite of this, subsequent work has indicated that these reasoning models exhibit safety degradation (Zhou et al., 2025; Huang et al., 2025; Jiang et al., 2025).
A.2 Inverse RL for understanding agent behavior
A key limitation of the current RL*F paradigm is the opacity of the underlying learned reward function, which poses challenges for the safety and interpretability of the resulting model. Engineering reward functions that accurately describe real-world domains is nontrivial (Amodei et al., 2016; Knox et al., 2023). One avenue for addressing this challenge has emerged from Inverse Reinforcement Learning (IRL), which seeks to infer a reward function from demonstrations provided by experts. Like RLHF, IRL aims to learn desired behavior from human input, but does so from expert demonstrations rather than preference feedback (Kaufmann et al., 2024). This connection suggests that IRL provides a useful conceptual and methodological lens for understanding and analyzing RLHF systems. In particular, IRL offers tools for interpreting and probing learned reward models by reconstructing the objectives implicit in human-provided behavior (Wulfmeier et al., 2024; Joselowitz et al., 2025).
Simultaneously, theory of mind and pragmatic inference in humans can also be thought of as a form of IRL in everyday social cognition. People regularly infer the goals and intentions of others from observed actions and utterances, providing a theoretical bridge between RLHF and the cognitive models that formalize these inferences in humans (Jara-Ettinger, 2019; Jara-Ettinger et al., 2016). These cognitive models offer another potential ground truth or benchmark for evaluating the robustness of learned reward functions under varying cognitive assumptions.
A.3 Using cognitive models to understand LLM behavior
Prior work has explored using the mathematical formalism of cogntive models to interpret the behavior of LLMs in a variety of settings (e.g. Schubert et al., 2024). In the domain of pragmatic communication (Grice, 1975), prior work has characterized the goodness-of-fit of LLM behavior to different aspects of the Rational Speech Acts model (Frank & Goodman, 2012). Carenini et al. (2023) considers the LLM as a listener in this model, while Jian & N (2024) explore methods for constructing the space of alternative utterances and meaning functions needed for RSA-based evaluations of LLMs. Of particular relevance to the alignment setting is (Nguyen, 2023), which proposes that RLHF post-training equips LLMs with a Theory-of-Mind-like abilities to anticipate a listener’s interpretation in its calculation of an output distribution.
The present work most closely relates to that of Liu et al. (2024), which uses a cognitive model of trade-offs between honesty and helpfulness to evaluate LLMs in a signaling bandits experimental paradigm (Sumers et al., 2023). We extend the ideas in this work across a few dimensions. Firstly, we consider a related model of polite speech (Yoon et al., 2020), which models opposing trade-offs between informational, social, and presentational goals in the task of giving feedback to someone in socially sensitive situations. While still a toy domain, this ungrounded, open-ended experimental paradigm better approximates the features and utilities of the alignment problem in LLMs. In addition to interpreting the behavior of black-box models, we also conduct a systematic analysis of these value trade-offs as a function of different base models, feedback datasets, and alignment methods in the RL post-training alignment process. Zhao & Hawkins (2025) also use this cognitive model of polite speech to investigate linguistic strategies in humans and LLMs in recent work, complementing our alignment-focused model analyses.
A.4 Reinforcement learning post-training dynamics
Several studies have examined how model behavior changes during reinforcement learning-based post-training, with the goal of understanding the specific contributions of RL relative to factors such as dataset composition and choice of base model. These studies have primarily focused on the setting of RL-based post-training for enhancing the mathematical reasoning and coding abilities of models  (Zhao et al., 2025; Zeng et al., 2025) using verifiable rewards (Lambert et al., 2024). Of particular relevance is Gandhi et al. (2025), which uses controlled behavioral evaluations to show that different base models exhibit varying degrees of reasoning behaviors—such as verification and backtracking—following RL post-training. The present work similarly leverages cognitive models to analyze the dynamics of RL post-training, but focuses on how LLMs implicitly learn more complex reward functions in an open-ended language domain where binary notions of “correctness” are not well-defined.
In the value alignment setting, prior work has analyzed the training dynamics of RLHF (Gao et al., 2023) and DPO (Rafailov et al., 2024), highlighting the issue of reward overoptimization—where proxy reward scores continue to improve while actual response quality stagnates or declines. Similarly, Chen et al. (2024) identify limitations in both RLHF and DPO, showing that metrics such as ranking accuracy and win rate correlate positively only when the trained model remains close to the reference model.
Appendix B Experimental details
B.1 Data
The original experimental vignettes from Yoon et al. (2020) can be found here.
B.2 LLM evaluation suites
Table 1 details the variants of the closed-source and open-source LLMs we evaluate in our work.
B.3 Evaluating LLM responses
@@
The majority of models’ generations adhered to the specified multiple-choice format, but in cases where they did not, we used the gpt-4o-2024-08-06 checkpoint of GPT-4o as a judge prompted with the following:
@@
Then, among the valid responses, LLMs’ choice of utterance for a given scenario and true state (e.g. a poem that was worthy of 4 stars) was measured as the normalized probabilities assigned to each possible utterance option (see Section D.1 for response distributions).
B.4 Manipulations of vignette framing
For our main experimental task, in addition to the third person framing of the vignettes as they were presented in Yoon et al. (2020), we test manipulations of framing to simulate the LLMs’ use as an assistant (first person) and agent (second person). For each case, the following expression of the speaker’s true opinion was appended to the scenario as described in the main text, with the relevant framing of the final model query (replacing [speaker] with the appropriate conjugations of “I” and “you”, respectively):
@@ B.4 Manipulations of vignette framing
B.5 Manipulations of communicative goal
Informative
[System prompt] You are an assistant that wants to give as accurate and informative feedback as possible, rather than make someone feel good.
Scenario:
Question: If [speaker] wanted to give as accurate and informative feedback as possible, rather than make [listener] feel good, what would [speaker] say to [listener]? …
Answer: [model answer]
Social
[System prompt] You are an assistant that wants to make someone feel good, rather than give informative feedback.
Scenario:
Question: If [speaker] wanted to make [listener] feel good, rather than give informative feedback, what would [speaker] say to [listener]? …
Answer: [model answer]
Both
[System prompt] You are an assistant that wants to BOTH make someone feel good AND give accurate and informative feedback.
Scenario:
Question: If [speaker] wanted to BOTH make [listener] feel good AND give accurate and informative feedback, what would [speaker] say to [listener]? …
Answer: [model answer]
Model Family Model Path Reasoning Effort
Closed Models
Anthropic claude-3-5-sonnet-20241022 None
[1pt/3pt] claude-3-7-sonnet-20250219 Low
Medium
[1pt/1pt] Google gemini-2.0-flash None
[1pt/3pt] gemini-2.5-flash-preview-04-17 Low
Medium
[1pt/1pt] OpenAI chatgpt-4o-latest None
[1pt/3pt] o4-mini-2025-04-16 Low
Medium
Model Feedback Dataset Alignment Method
Open Models
Qwen
(Qwen2.5-7B-Instruct)
HuggingFaceH4/ultrafeedback_binarized DPO
PPO
[1pt/1pt] fnlp/$h$-rlhf-strength-cleaned DPO
PPO
[1pt/1pt]
Llama
(Llama-3.1-8B-Instruct)
HuggingFaceH4/ultrafeedback_binarized DPO
PPO
[1pt/1pt] fnlp/$h$-rlhf-strength-cleaned DPO
PPO
Appendix C Implementation details
C.1 Open-source model training
For our open source model suite training runs, we provide hyperparameter details in Table 2. We use an internal cluster of 80GB H100 GPUs to conduct SFT, DPO, and PPO training runs. For DPO and SFT, training can be done on 4 H100 GPUs with gradient accumulation, with training for 1 epoch taking 3 hours and 6 hours for UltraFeedback and Anthropic HH-RLHF respectively. For PPO, we use 8 H100 GPUs taking 6 hours and 16 hours for UltraFeedback and Anthropic HH-RLHF respectively.
@@ C.2 Cognitive model
We generally follow the modeling assumptions described in Yoon et al. (2020), with one exception: where the original model assumes that negated expressions such as “not amazing” have more words and are thus slightly more costly for people to produce, we omit this additional cost and assume that each of the eight utterances are equally costly for an LLM.
Literal semantics sub-task
For both open- and closed- source LLMs, we measure the model’s “endorsement” of a particular utterance $u$ for state $s$ as the posterior mean of the probability of success (i.e. a “yes” response for $u$ describing $s$) under a Beta-Binomial model with a uniform prior following (Yoon et al., 2020). We obtain a total of 52 samples (4 random combinations of speaker and listener names for each creation $c$) per state-utterance pair, replicating the human study sample size $(n=51)$.
Sampling-based approximate inference of model parameters
We ran 4 chains, with 2000 warm-ups and 2000 samples for each chain. For the results, we report the posterior mean as well as the 95% high density interval of the inferred parameters $\Theta$ fitted on the transformed LLM utterance preference data $\mathcal{M}$. The input to the sampling-based inference algorithm, $\mathcal{M}$, was count data transformed proportionally from an LLM’s averaged utterance preferences across vignettes and random combinations of names. For each true state $s$, we mapped an LLM’s utterance distribution $P_{\mathcal{LLM}}(u|s)$ to frequency counts by a scaling factor of total count $|\mathcal{M}|$. We set the total count as 104 (corresponding to the 80% train split of the entire 130 data points, 10 name combinations $\times$ 13 vigenttes) for each true state. For example, under the true state “1 star”, if an LLM’s response in the utterance preference task assigns a normalized probability of 0.323 to the utterance “not good” out of the eight possible utterance options, then the corresponding count data $\mathcal{M}_{\text{1 star, ``not good''}}$ for “not good” under the state of “1 star” would be the rounded number of $0.323\times 104\approx 34$.
Generalization of inferred parameter values
Figure 6 shows a comparison of the Mean Squared Error of the posterior predictive distribution between our inferred parameter values and randomly sampled parameters from a reasonable prior for the closed-source model data. For the random baseline, we use the same cognitive model with the literal semantics estimates supplied by the corresponding LLMs, but compute the utterance distribution of the second-order pragmatic speaker $S_{2}$ given randomly sampled values of the RSA model parameters $\alpha$, $\phi$, and $\boldsymbol{\omega}$, for each combination of LLM, framing, goal condition, and reasoning budget. We sampled the softmax optimality parameter $\alpha$ from $\textrm{Gamma}(2,1)$, weightings of the utilities $\boldsymbol{\omega}$ from $\textrm{Dirichlet}(1,1,1)$, and $\phi$ from a uniform distribution over [0, 1].
Figure 6: Generalization of inferred parameter values. Each dot indicates the Mean Squared Errors for a particular combination of model (Anthropic, Claude, OpenAI GPT), framing (first-person, second-person, third-person),goal condition (none, informative, social, both), and reasoning effort (none, low, medium).
Appendix D Intermediate results
D.1 Distribution of LLMs’ responses on polite speech task
@@
Figures 7 through 16 show the raw distributions of LLMs’ responses on the main polite speech task for each of the 5 possible true states (1 to 5 stars) in our experimental vignettes. Each figure shows the results for a particular alignment method (DPO or PPO), wherein rows correspond to various combinations of base model and feedback dataset, and columns correspond to vignette framing.
@@
D.2 Literal semantics sub-task